Change the following point from rectangular to polar coordinate


‐ Square root of 3,2



Ii.change the following point from polar to rectangular coordinate



(3/2, n/12)


1
Expert's answer
2021-12-06T13:20:54-0500

1)

"r=\sqrt{\sqrt{3}^2+2^2}=\sqrt{7}\\
\phi=atan(\frac{2}{\sqrt{3}})\\(\sqrt{3},2)\rarr(\sqrt{7},atan(\frac{2}{\sqrt{3}}))"

2)

"x=r\cdot cos(\phi)=\frac{3}{2}\cdot cos(\frac{\pi}{12})"

"y=r\cdot sin(\phi)=\frac{3}{2}\cdot sin(\frac{\pi}{12})"

"(\frac{3}{2},\frac{\pi}{12})\rarr (\frac{3}{2}\cdot cos(\frac{\pi}{12}),\frac{3}{2}\cdot sin(\frac{\pi}{12}))"